using gordon growth model explain how covid 19 impact stock market negatively
Answer #1

Solution:

The Gordon growth model uses dividends, growth, and cost of equity to calculate the share price.

P0 = \frac{D_{1}}{(1+r)^1} + \frac{D_{2}}{(1+r)^2} + \frac{D_{3}}{(1+r)^3} + ......+ \frac{D_{n} + P}{(1+r)^n}

Where D1 = Dividend in year 1

D2 = Dividend in year 2

r = Cost of equity

Now due to covid 19 situation, the growth rate of the dividend will likely to go down. This lower dividend means that the share price will be low. Hence we can say that due to covid situation the price of the stocks will be lower.